# 3]Sorbonne Université, CNRS, Université de Paris, Inria, Laboratoire Jacques-Louis Lions UMR 7598, 75005 Paris, France

### v1

@@ 4.2.1 Travelling waves in PDE models of spatial spread dynamics
so that the stable homogeneous steady-state “invades” the unstable one. The boundary conditions (26) then corresponds to considering a scenario in which the population spreads across the uncolonised surrounding space. A classical result is that there exist monotonically decreasing solutions of the problem (25)-(26), provided that $c\geq c^{*}:=2\ \sqrt{rD}$, where the threshold value of the speed $c^{*}$ is called the minimal wave speed [191, 204].
@@
converges (in some appropriate sense) to $\rho^{0}(x-c^{*}t)$, with $c^{*}=2$ [89, 105]. A corroboration of this result is shown in Figure 3, where numerical solutions of (28)-(29) are seen to become increasingly step-like as $\varepsilon\rightarrow 0$ and with a numerically calculated wave speed $v\approxeq 2$. As shown in [33, 108, 89], a useful tool to prove this and related asymptotic results is the real phase WKB ansatz borrowed from geometric optics, which is also referred to as the Hopf-Cole transformation:
@@
A key observation underlying the change of variable (30) is that if $u_{\varepsilon}(t,x)$ converges (uniformly, locally in time) to $u(t,x)$ as $\varepsilon\to 0$, then $\rho(t,x)=0$ and $\rho(t,x)=1$ correspond to $u(t,x)<0$ and $u(t,x)=0$, respectively. Hence, the limiting behaviour of $\rho_{\varepsilon}(t,x)$ can be characterised by studying the limiting behaviour of $u_{\varepsilon}(t,x)$. The advantage of this approach lies in the fact that upon substituting the ansatz (30) into the rescaled non-local PDE (28), one finds that $u_{\varepsilon}(t,x)$ satisfies a type of Hamilton-Jacobi equation for which there is a range of mathematical tools, stemming from the work of Lions and coworkers in the 1980s on viscosity solutions of Hamilton-Jacobi equations [58, 50, 162]. This in turn facilitates studies into the behaviour of $\rho_{\varepsilon}(t,x)$ as $\varepsilon\to 0$.
@@ 4.2.2 Concentration phenomena in non-local PDE models of evolutionary dynamics
Here, the parameters $\rho^{0}$ and $\bar{y}^{0}$ model the initial size of the population and the mean or prevailing phenotype in the population at time $t=0$, respectively. The parameter $\sigma_{0}^{2}$ is the related variance, which provides a measure of intra-population phenotypic heterogeneity. It is possible to prove (see, for instance,  Almeida et al. 1, Ardaševa et al. 9, Chisholm et al. 61) that, when subject to the initial condition (31), the non-local PDE model (10)-(12) with $y\in\mathbb{R}$ admits solutions of the Gaussian form
@@
From (33) and (32), one sees that, when subject to the initial condition (37), the rescaled non-local PDE obtained by complementing (35) with (11)-(12) admits the solution
@@
$\rho(t)=\dfrac{\gamma-\left(\bar{y}(t)-\varphi\right)^{2}}{k}\;\;\text{for a.e.}\;\;t\in(0,\infty)\,.$ (42)
This is also confirmed by the results of numerical simulations of (35), subject to (37) and complemented with (11)-(12), which are displayed in Figure 5. Here, numerical solutions are seen to be of the Gaussian form (38), which becomes concentrated as a weighted infinitely sharp Gaussian (i.e. a weighted Dirac mass) in the asymptotic regime $\varepsilon\rightarrow 0$.
@@
Moreover, as proposed in [75], the concentration point (i.e. the centre of the Dirac mass) $\bar{y}(t)$ can be biologically interpreted as the trait that is shown by the population at time $t$. Hence, the ODE (41)1 can be regarded as a type of canonical equation of adaptive dynamics: an ODE that describes how the prevailing trait in the population changes over time [181]. From the Cauchy problem (41) it is easy to see that $\bar{y}(t)\to\varphi$ as $t\to\infty$. Recalling that the parameter $\varphi$ models the fittest phenotype (see the fitness function defined via (11)-(12)), the latter convergence result translates into mathematical terms the concept of “survival of the fittest” originating from evolutionary theory [222].
Notably, the solution to the rescaled non-local PDE (35) also exhibits concentration phenomena for a broader class of fitness functions and initial conditions, provided they have essentially the same structural properties of the fitness function defined via (11)-(12) and the initial condition (37) – in other words, functions $R(y,\rho)$ that are strictly concave in $y$ and strictly monotonically decreasing in $\rho$, and initial conditions $n^{0}_{\varepsilon}(y)$ that converge (in the weak sense of measures) to weighted Dirac masses as $\varepsilon\to 0$.
A robust method for studying such concentration phenomena was originally proposed in [75, 202], and then developed and extended in [41, 37, 63, 153, 174, 167, 179, 200]. This built on the observation that, given that the rescaled non-local PDE (35)1 is of close form to the rescaled Fisher-KPP equation (28), with the exception of how the reaction term is defined, it is natural to make a change of variable of the same type as (30), that is, use the WKB ansatz
@@
$n^{0}_{\varepsilon}(y)\xrightharpoonup[\varepsilon\rightarrow 0]{}\rho^{0}\delta_{\bar{y}^{0}}(y)\,,\quad\rho_{0}\in\mathbb{R}^{+}_{*}\,,\;\;\bar{y}_{0}\in\mathbb{R}\,.$
@@
Here $u(t,y)$ – which is the limit of $u_{\varepsilon}(t,y)$ as $\varepsilon\to 0$ – is a viscosity solution (in the sense introduced in [41, 37, 200]) of the constrained Hamilton-Jacobi equation
@@
As final remarks, the aforementioned results on Gaussian solutions and concentration phenomena extend to the case where the fitness function is also a periodic function of $t$, which corresponds to the situation of populations exposed to periodically fluctuating environments, as demonstrated, for instance, in [9, 156] and [96, 188]. Furthermore, concentration phenomena have been investigated also in non-local PDE models of evolutionary dynamics of forms similar to (10) but where the diffusion term is replaced by an advection term modelling cell differentiation, with an advection velocity that is a function of the phenotype structure [118].
4.2.3 Travelling waves and concentration phenomena in PS-PIDE models of spatial and evolutionary dynamics
@@
In the context of PS-PIDE models of type (2), travelling waves are solutions of the form
@@
which generalises the form (24) to the case where a phenotypic structure is incorporated into the model. Building on the formal results presented in [28], particular attention has been given to travelling-wave solutions of the PS-PIDE model (8)-(9) with ${\bm{x}}\equiv x\in\mathbb{R}$ and ${\bm{y}}\equiv y\in(0,Y)$, where $Y\in\mathbb{R}^{+}_{*}$, and subject to zero-flux boundary conditions at $y=0$ and $y=Y$.
Summarising, the existence of solutions to the boundary value problem defined by complementing the PDE obtained by substituting (47) into the PS-PIDE model (8)-(9) with appropriate boundary conditions at $z=-\infty$ and $z=\infty$ was proven in [26], where the corresponding minimal wave speed was also characterised. The convergence of the solution to the PS-PIDE model (8)-(9) to travelling waves, for large $t$ and $x$ and under appropriate assumptions on the initial condition, was subsequently studied in [233]. The first step of the study carried out in [233] employs the space and time scaling (27), which leads to the following rescaled PS-PIDE model
@@
into (48)1. Note that (49) is a natural extension of (30) to the case where the solution to the model equation depends both on the spatial variable $x$ and the phenotypic variable $y$. Convergence of the solution to the PS-PIDE model (8)-(9) to travelling waves as $t\to\infty$, again under appropriate assumptions on the initial condition, was also studied in [38, 35], where a detailed characterisation of the dynamics of the maximum point of the phenotype density function $n$ along the phenotypic dimension $y$ (i.e. the dynamics of the prevailing phenotype) at the edge of the wave front was also established. The emergence of accelerating fronts when $Y=\infty$ (i.e. in the case of unbounded motility) was also investigated in [38, 28, 35].
@@
In [238], it was proven that the PS-PIDE model (13)-(15) with ${\bm{x}}\equiv x\in\mathcal{X}\subset\mathbb{R}$, ${\bm{y}}\equiv y\in\mathbb{R}$, $D=0$ (i.e. cell movement is neglected), and ${\bm{S}}(t,{\bm{x}})\equiv{\bm{\Sigma}}(x)$, where the ${\bm{\Sigma}}(x)$ is given (i.e. the spatially heterogeneous environment in which the population is embedded does not evolve in time), and subject to initial condition (50), admits solutions of the Gaussian form
@@
In fact, concentration phenomena have also been studied for rescaled PS-PIDE models of form
@@
which can be obtained from PS-PIDE models like (13), when ${\bm{x}}\equiv x\in\mathcal{X}$, ${\bm{y}}\equiv y\in\mathcal{Y}$, $D({\bm{y}})\equiv D$, and ${\bm{S}}(t,{\bm{x}})\equiv S(t,x)$, by employing the space and time scaling (27) along with the parameter scaling (34).
To provide greater detail, using the WKB ansatz (49) – a natural choice in view of the similarities between the rescaled PS-PIDE (52)1 and the rescaled non-local PDE (35)1 – and making appropriate assumptions on the initial condition, $n_{\varepsilon}(0,x,y)$, the fitness function, $R$, and the governing equation for $S_{\varepsilon}$, asymptotic results can be obtained that express concentration phenomena analogous to (40), i.e.
@@
These were proven in [139], building on the results presented in [186], taking $D=0$ and $\mathcal{Y}\equiv\mathbb{R}$. In addition, when $\mathcal{X}\equiv\mathbb{R}$, assuming
$n^{0}_{\varepsilon}(y)\xrightharpoonup[\varepsilon\rightarrow 0]{}\rho^{0}(x)\delta_{\bar{y}^{0}(x)}(y)\ ,\quad\rho^{0}:\mathbb{R}\to\mathbb{R}^{+}_{*}\ ,\quad\bar{y}_{0}:\mathbb{R}\to\mathbb{R}\ ,$
one can show, as performed in [237] taking $D>0$, that the concentration point $\bar{y}(t,x)$ in (53) formally satisfies the following Cauchy problem
@@
and it is such that $\partial^{2}_{yy}u(t,x,\bar{y}(t,x))<0$. Moreover, if the function $R(y,\rho,S)$ is strictly monotonically decreasing in $\rho$, the weight $\rho(t,x)>0$ in (53) is given as a function of $\bar{y}(t,x)$ and $S(t,x)$ by the relation
@@
As a concluding remark, we note that concentration phenomena of type (53) have also been investigated in PS-PIDE models of form (13) with $D({\bm{y}})\equiv D$ (see Bouin and Mirrahimi 36) and in more general PS-PIDE models of related forms (see Hao et al. 127).
@@
More recently, concentration phenomena have been investigated across travelling waves in PS-PIDE models for the spatial spread and evolutionary dynamics of cell populations under the pressure-based form (17) and the taxis-based form (20), with ${\bm{x}}\equiv x\in\mathbb{R}$ and ${\bm{y}}\equiv y\in(0,Y)$ where $Y\in\mathbb{R}^{+}_{*}$. Applying the space and time scaling (27) and the parameter scaling (34), rescaled models of the type of (20) are given by
@@
Both rescaled PS-PIDEs are subject to zero-flux boundary conditions at $y=0$ and $y=Y$.
In particular, under appropriate assumptions on the model functions, including that the function $R$ is strictly monotonically decreasing in $\rho$, employing the ansatz (49), it was shown that rescaled PS-PIDEs of form (56) (see Lorenzi and Painter 172, Lorenzi et al. 168) and rescaled PS-PIDEs of form (57) (see Lorenzi et al. 173, Macfarlane et al. 189) formally admit travelling-wave solutions $n_{\varepsilon}(t,x,y)\equiv n_{\varepsilon}(z,y)$, with $c\in\mathbb{R}^{+}_{*}$, such that
@@
These correspond to biological scenarios wherein the cell population is monomorphic at each position along the wave, and the concentration point $\bar{y}(z)$ and the weight $\rho(z)$ represent, respectively, the trait that is expressed by the cells and the cell density at position $z$. For the rescaled PS-PIDE model (56), $\bar{y}(z)$ formally satisfies the ODE
@@
The ODE (59) and the relation (60) are coupled with an ODE for $S(z)$, which is the travelling-wave solution of the equation for $S_{\varepsilon}(t,x)$ as $\varepsilon\to 0$. Analogously, for the rescaled PS-PIDE model (57), $\bar{y}(z)$ formally satisfies the ODE
@@
which are common assumptions to make. The function $u(z,y)$ in (59) (or (61)) formally satisfies a Hamilton-Jacobi equation subject to the constraint
@@
Note that assumptions (63) ensure that $\Pi^{-1}(P_{M})$ is the maximum value of $\rho(z)=\Pi^{-1}[P](z)$ if $0\leq P\leq P_{M}$. Under the boundary condition (66), through direct calculations on (59)-(60) and (61)-(62), it is then possible to show that, in both cases, there is a minimal wave speed $c^{*}\in\mathbb{R}^{+}_{*}$ such that if $c>c^{*}$ then the concentration point $\bar{y}(z)$ and the weight $\rho(z)$ in (58) formally satisfy the following relations
@@
with ${}\ell\in\mathbb{R}\cup\{\infty\}$. Under the aforementioned monotonicity assumptions on the functions $r(y)$, $\chi(y)$, and $\mu(y)$, these results provide a mathematical formalisation of the idea that cells with a more migratory and less proliferative phenotype (i.e. in phenotypic states corresponding to $y$ close to $Y$) are concentrated towards the front of the travelling wave, whereas cells with a less migratory and more proliferative phenotype (i.e. in phenotypic states corresponding to $y$ close to 00) make up the population bulk in the rear of the wave. Numerical simulations of the rescaled pressure-based PS-PIDE model (57) and the rescaled taxis-based PS-PIDE model (56) show this within-wave phenotype structuring (see Figure 6(b), left panel and Figure 7(a), left panel), where we also note the verification of the minimal wave speed (see Figure 6(b), right panel and Figure 7(a), right panel).
4.3 Tools and techniques to simulate PS-PIDE models
When simulating PS-PIDEs of form (2), numerical challenges may arise due to the potential composite shapes and sharp features of the solution, or the stiff [^16] A stiff PDE generally contains terms that can lead to rapid variations in the solution, to the extent that methods for numerical integration in time would generally require extremely small time-steps to ensure stability. This concept naturally extends to PDE systems, where dynamics modelled by different equations may occur over different timescales. nature of the problems. We offer a brief step-by-step guide into the techniques that can be used to overcome these challenges, when employing numerical schemes that rely on the method of lines (MOL) [134]. We illustrate through the following one-dimensional, in each of physical and phenotype space, PS-PIDE model

### v2

@@ 4.2.1 Travelling waves in PDE models of spatial spread dynamics
so that the stable homogeneous steady-state “invades” the unstable one. The boundary conditions (26) then correspond to considering a scenario in which the population spreads across the uncolonised surrounding space. A classical result is that there exist monotonically decreasing solutions of the problem (25)-(26), provided that $c\geq c^{*}:=2\ \sqrt{rD}$, where the threshold value of the speed $c^{*}$ is called the minimal wave speed [195, 207].
@@
converges (in some appropriate sense) to $\rho^{0}(x-c^{*}t)$, with $c^{*}=2$ [91, 108]. A corroboration of this result is shown in Figure 3, where numerical solutions of (28)-(29) are seen to become increasingly step-like as $\varepsilon\rightarrow 0$ and with a numerically calculated wave speed $v\approxeq 2$. As shown by [22, 101, 91], a useful tool to prove this and related asymptotic results is the real phase WKB ansatz borrowed from geometric optics, which is also referred to as the Hopf-Cole transformation:
@@
A key observation underlying the change of variable (30) is that if $u_{\varepsilon}(t,x)$ converges (uniformly, locally in time) to $u(t,x)$ as $\varepsilon\to 0$, then $\rho(t,x)=0$ and $\rho(t,x)=1$ correspond to $u(t,x)<0$ and $u(t,x)=0$, respectively. Hence, the limiting behaviour of $\rho_{\varepsilon}(t,x)$ can be characterised by studying the limiting behaviour of $u_{\varepsilon}(t,x)$. The advantage of this approach lies in the fact that upon substituting the ansatz (30) into the rescaled (local) PDE (28), one finds that $u_{\varepsilon}(t,x)$ satisfies a type of Hamilton-Jacobi equation for which there is a range of mathematical tools, stemming from the work of Lions and coworkers in the 1980s on viscosity solutions of Hamilton-Jacobi equations [66, 67, 162]. This in turn facilitates studies into the behaviour of $\rho_{\varepsilon}(t,x)$ as $\varepsilon\to 0$.
@@ 4.2.2 Concentration phenomena in non-local PDE models of evolutionary dynamics
Here, the parameters $\rho^{0}$ and $\bar{y}^{0}$ model the initial size of the population and the mean or prevailing phenotype in the population at time $t=0$, respectively. The parameter $\sigma_{0}^{2}$ is the related variance, which provides a measure of intra-population phenotypic heterogeneity. It is possible to prove (see, for instance,  Almeida et al 9, Ardaševa et al 15, Chisholm et al 61) that, when subject to the initial condition (31), the non-local PDE model (10)-(12) with $y\in\mathbb{R}$ admits solutions of the Gaussian form
@@
From (32) and (33), one sees that, when subject to the initial condition (37), the rescaled non-local PDE obtained by complementing (35) with (11),(12) admits the solution
@@
$\rho(t)=\dfrac{\gamma-\left(\bar{y}(t)-\varphi\right)^{2}}{\kappa}\;\;\text{for a.e.}\;\;t\in(0,\infty)\,.$ (42)
This is also confirmed by the results of numerical simulations of (35), subject to (37) and complemented with (11),(12), which are displayed in Figure 5. Here, numerical solutions are seen to be of the Gaussian form (38), which becomes concentrated as a weighted infinitely sharp Gaussian (i.e. a weighted Dirac mass) in the asymptotic regime $\varepsilon\rightarrow 0$.
@@
Moreover, as proposed by [78], the concentration point (i.e. the centre of the Dirac mass) $\bar{y}(t)$ can be biologically interpreted as the trait that is shown by the population at time $t$. Hence, the ODE (41)1 can be regarded as a type of canonical equation of adaptive dynamics: an ODE that describes how the prevailing trait in the population changes over time [188]. From the Cauchy problem (41) it is easy to see that $\bar{y}(t)\to\varphi$ as $t\to\infty$. Recalling that the parameter $\varphi$ models the fittest phenotype (see the fitness function defined via (11),(12)), the latter convergence result translates into mathematical terms the concept of “survival of the fittest” originating from evolutionary theory [227].
Notably, the solution to the rescaled non-local PDE (35) also exhibits concentration phenomena for a broader class of fitness functions and initial conditions, provided they have essentially the same structural properties of the fitness function defined via (11),(12) and the initial condition (37) – in other words, functions $R(y,\rho)$ that are strictly concave in $y$ and strictly monotonically decreasing in $\rho$, and initial conditions $n^{0}_{\varepsilon}(y)$ that converge (in the weak sense of measures) to weighted Dirac masses as $\varepsilon\to 0$.
A robust method for studying such concentration phenomena was originally proposed by [78, 205], and then developed and extended by [21, 23, 62, 155, 177, 174, 193, 208]. This built on the observation that, given that the rescaled non-local PDE (35)1 is of close form to the rescaled Fisher-KPP equation (28), with the exception of how the reaction term is defined, it is natural to make a change of variable of the same type as (30), that is, use the WKB ansatz
@@
$n^{0}_{\varepsilon}(y)\xrightharpoonup[\varepsilon\rightarrow 0]{}\rho^{0}\delta_{\bar{y}^{0}}(y)\quad\text{(weakly in measures)}\,,\quad\rho_{0}\in\mathbb{R}^{+}_{*}\,,\;\;\bar{y}_{0}\in\mathbb{R}\,.$
@@
Here $u(t,y)$ – which is the limit of $u_{\varepsilon}(t,y)$ as $\varepsilon\to 0$ – is a viscosity solution (in the sense introduced by [21, 23, 208]) of the constrained Hamilton-Jacobi equation
@@
As final remarks, the aforementioned results on Gaussian solutions and concentration phenomena extend to the case where the fitness function is also a periodic function of $t$, which corresponds to the situation of populations exposed to periodically fluctuating environments, as demonstrated, for instance, by [15, 167] and [97, 194]. Furthermore, concentration phenomena have been investigated also in non-local PDE models of evolutionary dynamics of forms similar to (10) but where the diffusion term is replaced by an advection term modelling cell differentiation, with an advection velocity that is a function of the phenotype structure [119].
4.2.3 Travelling waves and concentration phenomena in PS-PDE models of spatial and evolutionary dynamics
@@
In the context of PS-PDE models of type (2), travelling waves are solutions of the form
@@
which generalises the form (24) to the case where a phenotypic structure is incorporated into the model. Building on the formal results presented by [34], particular attention has been given to travelling-wave solutions of the PS-PDE model (8)-(9) with ${\bm{x}}\equiv x\in\mathbb{R}$ and ${\bm{y}}\equiv y\in(0,Y)$, where $Y\in\mathbb{R}^{+}_{*}$, and subject to zero-flux boundary conditions at $y=0$ and $y=Y$.
Summarising, the existence of solutions to the boundary value problem defined by complementing the PDE obtained by substituting (47) into the PS-PDE model (8)-(9) with appropriate boundary conditions at $z=-\infty$ and $z=\infty$ was proven by [32], who also characterised the corresponding minimal wave speed. The convergence of the solution to the PS-PDE model (8)-(9) to travelling waves, for large $t$ and $x$ and under appropriate assumptions on the initial condition, was subsequently studied by [238]. The first step of the study carried out by [238] employs the space and time scaling (27), which leads to the following rescaled PS-PDE model
@@
into (48)1. Note that (49) is a natural extension of (30) to the case where the solution to the model equation depends both on the spatial variable $x$ and the phenotypic variable $y$. Convergence of the solution to the PS-PDE model (8)-(9) to travelling waves as $t\to\infty$, again under appropriate assumptions on the initial condition, was also studied by [29, 35], who established a detailed characterisation of the dynamics of the maximum point of the phenotype density function $n$ along the phenotypic dimension $y$ (i.e. the dynamics of the prevailing phenotype) at the edge of the wave front as well. The emergence of accelerating fronts when $Y=\infty$, which, for the choice of $D(y)$ given by (9) (i.e. $D(y):=y$), corresponds to the case of unbounded motility, was also investigated by [29, 34, 35].
@@
[244] proved that the PS-PDE model (13)-(15) with ${\bm{x}}\equiv x\in\mathcal{X}\subset\mathbb{R}$, ${\bm{y}}\equiv y\in\mathbb{R}$, $D(y)\equiv 0$ (i.e. cell movement is neglected), and ${\bm{S}}(t,{\bm{x}})\equiv{\bm{\Sigma}}(x)$, where ${\bm{\Sigma}}(x)$ is given (i.e. the spatially heterogeneous environment in which the population is embedded does not evolve in time), and subject to initial condition (50), admits solutions of the Gaussian form
@@
In fact, concentration phenomena have also been studied for rescaled PS-PDE models of form
@@
which can be obtained from PS-PDE models like (13), when ${\bm{x}}\equiv x\in\mathcal{X}$, ${\bm{y}}\equiv y\in\mathcal{Y}$, $D({\bm{y}})\equiv D$, and ${\bm{S}}(t,{\bm{x}})\equiv S(t,x)$, by employing the space and time scaling (27) alongside the parameter scaling (34).
To provide greater detail, using the WKB ansatz (49) – a natural choice in view of the similarities between the rescaled PS-PDE (52)1 and the rescaled non-local PDE (35)1 – and making appropriate assumptions on the initial condition, $n_{\varepsilon}(0,x,y)$, the fitness function, $R$, and the governing equation for $S_{\varepsilon}$, asymptotic results can be obtained that express concentration phenomena analogous to (40), i.e.
@@
These were proven by [138], building on the results presented by [192] for the case where $D=0$ and $\mathcal{Y}\equiv\mathbb{R}$. In addition, when $\mathcal{X}\equiv\mathbb{R}$, assuming
$n^{0}_{\varepsilon}(y)\xrightharpoonup[\varepsilon\rightarrow 0]{}\rho^{0}(x)\delta_{\bar{y}^{0}(x)}(y)\quad\text{(weakly in measures)}\,,\quad\rho^{0}:\mathbb{R}\to\mathbb{R}^{+}_{*}\ ,\quad\bar{y}_{0}:\mathbb{R}\to\mathbb{R}\ ,$
one can formally show, as performed by [245] taking $D>0$, that the concentration point $\bar{y}(t,x)$ in (53) formally satisfies the following Cauchy problem
@@
and it is such that $\partial^{2}_{yy}u(t,x,\bar{y}(t,x))<0$. Moreover, if the function $R(y,\rho,S)$ is strictly monotonically decreasing in $\rho$, the weight $\rho(t,x)>0$ in (53) is given as a function of both $\bar{y}(t,x)$ and $S(t,x)$ by the relation
@@
As a concluding remark, we note that concentration phenomena of type (53) have also been investigated in PS-PDE models of form (13) with $D({\bm{y}})\equiv D$, see [33], and in more general PS-PDE models of related forms, see [127].
@@
More recently, concentration phenomena have been investigated across travelling waves in PS-PDE models for the spatial spread and evolutionary dynamics of cell populations under the pressure-based form (17) and the taxis-based form (20), with ${\bm{x}}\equiv x\in\mathbb{R}$ and ${\bm{y}}\equiv y\in(0,Y)$ where $Y\in\mathbb{R}^{+}_{*}$. Applying the space and time scaling (27) and the parameter scaling (34), alongside the scaling $\mu(y)\to\dfrac{\mu(y)}{\varepsilon}$ for (17) and the scaling $\chi(y)\to\dfrac{\chi(y)}{\varepsilon}$ for (20), rescaled models of the type of (20) are given by
@@
Both rescaled PS-PDEs are subject to zero-flux boundary conditions at $y=0$ and $y=Y$.
In particular, under appropriate assumptions on the model functions, including that the function $R$ is strictly monotonically decreasing in $\rho$, employing the ansatz (49), it was shown that rescaled PS-PDEs of form (56), see [165, 173], and rescaled PS-PDEs of form (57), see [171, 182], formally admit travelling-wave solutions $n_{\varepsilon}(t,x,y)\equiv n_{\varepsilon}(z,y)$, with $c\in\mathbb{R}^{+}_{*}$, such that
@@
These correspond to biological scenarios wherein the cell population is monomorphic at each position along the wave, and the concentration point $\bar{y}(z)$ and the weight $\rho(z)$ represent, respectively, the trait that is expressed by the cells and the cell density at position $z$. For the rescaled PS-PDE model (56), $\bar{y}(z)$ formally satisfies the ODE
@@
The ODE (59) and the relation (60) are coupled with an ODE for $S(z)$, which is the travelling-wave solution of the equation for $S_{\varepsilon}(t,x)$ as $\varepsilon\to 0$. Analogously, for the rescaled PS-PDE model (57), $\bar{y}(z)$ formally satisfies the ODE
@@
which are common assumptions to make [13]. The function $u(z,y)$ in (59) (or (61)) formally satisfies a Hamilton-Jacobi equation subject to the constraint
@@
Note that assumptions (63) ensure that if $0\leq P\leq P_{M}$ then $\Pi^{-1}(P_{M})$ is the maximum value of $\rho(z)=\Pi^{-1}[P](z)$. Under the boundary condition (66), through direct calculations on (59),(60) and (61),(62), it is possible to show that, in both cases, there is a minimal wave speed $c^{*}\in\mathbb{R}^{+}_{*}$ such that if $c>c^{*}$ then the concentration point $\bar{y}(z)$ and the weight $\rho(z)$ in (58) formally satisfy the following relations
@@
with ${}\ell\in\mathbb{R}\cup\{\infty\}$. Under the aforementioned monotonicity assumptions on the functions $r(y)$, $\chi(y)$, and $\mu(y)$, these results provide a mathematical formalisation of the idea that cells with a more migratory and less proliferative phenotype (i.e. in phenotypic states corresponding to $y$ close to $Y$) are concentrated towards the front of the travelling wave, whereas cells with a less migratory and more proliferative phenotype (i.e. in phenotypic states corresponding to $y$ close to 00) make up the population bulk in the rear of the wave. Numerical simulations of the rescaled pressure-based PS-PDE model (57) and the rescaled taxis-based PS-PDE model (56) show this within-wave phenotype structuring (see Figure 6(b), left panel and Figure 7(a), left panel), where we also note the verification of the minimal wave speed (see Figure 6(b), inset of the right panel and Figure 7(a), right panel).
4.3 Tools and techniques to simulate PS-PDE models
When simulating PS-PDEs of form (2), numerical challenges may arise due to the potential composite shapes and sharp features of the solution, or the stiff [^18] A stiff PDE generally contains terms that can lead to rapid variations in the solution, to the extent that methods for numerical integration in time would generally require extremely small time-steps to ensure stability. This concept naturally extends to PDE systems, where dynamics modelled by different equations may occur over different timescales. nature of the problems. We offer a brief step-by-step guide into the techniques that can be used to overcome these challenges, when employing numerical schemes that rely on the method of lines (MOL) [136]. We illustrate through the following one-dimensional, in each of physical and phenotype space, PS-PDE model